A container is filled to the brim with 1.5 L of mercury at 20°C. As the temperature of the container and mercury is increased to 56°C, a total of 7.8 mL of mercury spill over the brim of the container. Determine the linear expansion coefficient of the material that makes up the container.
Answer #1

The linear expansion coefficient of the container is one- third its coefficient of volume expansion.

αc = βc/3    or      αHg = βHg/3    -------------(1)      

After heating change in volume of the container = ∆Vc = βcVc∆T   -----------------(2)

After heating change in volume of the mercury = ∆VHg = βHgVHg∆T   -----------------(3)

∆VHg - ∆Vc = 7.8*10-3L --------------------(4)

From eqn (2), (3) and (4)

βHgVHg∆T   - βcVc∆T   = 7.8*10-3L

βc = (βHgVHg∆T   - 7.8*10-3L)/( Vc∆T)

but VHg = Vc = V

βc = (βHgV∆T   - 7.8*10-3L)/( V∆T)

βc = βHg    - (7.8*10-3L)/( V∆T)

From eqn (1),

αc = βHg/3   - (7.8*10-3L)/( 3V∆T)

Now plugging values βHg = (0.18*10-3K-1)

αc = (0.18*10-3K-1)/3   - (7.8*10-3L)/( 3(1.5)(36)) = 1.2*10-5 K-1